OpenStudy (anonymous):

Write the equation of the graph. ALGTests_L2_S1_U7_Q12v2Stem.gif A. y – 3 = ½(x + 1) B. y + 3 = –½(x – 1) C. y = ½x + 5 D. y – 3 = –½(x + 1)

OpenStudy (anonymous):

Copy and paste link into search bar @TheSmartOne

OpenStudy (anonymous):

@Utterly_Confuzzled

TheSmartOne (thesmartone):

That link won't work :P

OpenStudy (anonymous):

Here is the new one.

OpenStudy (utterly_confuzzled):

y=mx+b b is the y-intercept. Based on the graph you can see that the y-int is 5 correct? That becomes y=mb+5 Next, find the slope and replace m with it

OpenStudy (anonymous):

1/2x

OpenStudy (utterly_confuzzled):

Right. So the formula would become y=1/2x+5

OpenStudy (anonymous):

Okay thanks @TheSmartOne would that be correct?

TheSmartOne (thesmartone):

That's correct. Great job, @Abbster2015 and @Utterly_Confuzzled :)

OpenStudy (anonymous):

Thanks! Only one more :)